OpenStudy (anonymous):

Convert each subtraction problem to addition and type in the matching number. 3/4 - -2/3 =

OpenStudy (lalaly):

\[\frac{3}{4} + \frac{2}{3}\]

OpenStudy (anonymous):

3/4 + 2/3 = 9/12+ 8/12 = 17/12= 1 and 5/12